Ellie is 6 and her dad is 37. How long will it be before her dad is twice as old as she is?
Answer by hamsanash1981@gmail.com(151)

Let n be the year after her dad is twice old as her.
37+x = 2(6+x)
37+x = 12+2x
X = 25
It will take 25 years for her dad's age = 2 * her age.
It will be 24 years long before it happens so.